# MO-YOLO: End-to-End Multiple-Object Tracking Method with YOLO and Decoder

### v3

@@ Abstract
The recent emergence of Transformer-based end-to-end models, exemplified by MOTR, has showcased remarkable multi-object tracking (MOT) performance on datasets like DanceTracker. However, the computational complexities associated with these models pose challenges in both training and deployment. Drawing insights from successful models such as GPT, our proposed MO-YOLO stands out as an efficient and computationally frugal end-to-end MOT solution. MO-YOLO integrates principles from YOLO and RT-DETR, adopting a decoder-centric architecture alongside other complementary structures. By leveraging the RT-DETR decoder and architectural components from YOLOv8, MO-YOLO achieves high-speed performance, shorter training times, and proficient MOT capabilities. On the Dancetrack dataset, MO-YOLO not only achieves parity with MOTR in tracking performance but also surpasses it in terms of speed (MOTR 9.5 FPS, MO-YOLO 19.6 FPS). Moreover, MO-YOLO demonstrates significantly reduced training times and lower hardware requirements compared to MOTR. This research introduces a promising paradigm for efficient end-to-end MOT, highlighting enhanced performance and resource efficiency in a succinct and impactful manner.
Keywords:
Multiple-Object Tracking, Transformer ,YOLO, End-to-End
@@ 1 Introduction
Multi-Object Tracking (MOT) is a crucial task in computer vision, aiming to predict object trajectories across consecutive images [2]. The tracking-by-detection approach, relying on object detectors for individual target identification [1, 22, 9], has gained prominence. This method involves associating detections using algorithms like the Kalman filter [3] to form object trajectories. While demonstrating remarkable performance, the intrinsic two-step process—commencing with initial object detection and followed by object tracking—introduces heightened computational complexity and the potential for error propagation.
In recent times, the field has witnessed the emergence of end-to-end MOT models based on Transformer architectures [38, 40, 42], boasting superior performance on datasets such as DanceTracker [29] compared to tracking-by-detection methods. However, the computational demands associated with Transformer models present challenges in terms of both training and deployment, necessitating formidable hardware resources.
In order to address these challenges more effectively, a retrospective analysis of the origins of the Transformer model becomes imperative. The Transformer architecture was initially proposed for the domain of Natural Language Processing (NLP) [31]. As researchers continuously refined this model, the GPT[4] emerged as a prominent highlight in the NLP field over the past few years. GPT, exclusively utilizing the decoder component of the Transformer, achieves pre-training and fine-tuning for various language tasks by stacking multiple decoder layers to construct a deep neural network.
Inspired by the success of GPT, this paper adopts a network structure combining the principles of You Only Look Once (YOLO)[25] and RT-DETR[21], incorporating tracking concepts from MOTR[42]. The resultant model, named MO-YOLO, follows a decoder-based end-to-end MOT approach.
MO-YOLO leverages the decoder from RT-DETR and the block and neck components from YOLOv8 [14], offering advantages such as high speed and shorter training times. This architecture demonstrates proficient performance in multitarget tracking tasks.
This paper pioneers a fusion of concepts from YOLO and the end-to-end MOTR model, resulting in a comprehensive end-to-end MOT model. The contributions of this work can be encapsulated in four fundamental advancements:
@@
    The primary contribution involves the development of a novel end-to-end tracking network. This network amalgamates structural components derived from YOLO, RT-DETR, and the conceptual framework of MOTR, establishing a cohesive and efficient model for Multiple Object Tracking.
@@
    The third contribution introduces the Tracking Box Selection Process (TBSP) strategy, strategically filtering bounding boxes during training to expedite model convergence and enhance training process efficiency.
  • 4)
    Lastly, we substantiate the feasibility of a decoder based model akin to GPT in the domain of MOT in computer vision.
@@ 2 Related work
Figure 1: MOTR’s Streamline the process of updating detect queries and track queries under typical MOT scenarios. The track query set, initially empty, is dynamically adjusted in length. Detect queries play a role in identifying new objects.
@@ 2.1 MOTR Series
MOTR [42] is a groundbreaking model that achieves end-to-end multi-object tracking. It extends the concept of object queries from DETR [7] to track queries, allowing for dynamic tracking in video sequences. By employing a learnable positional embedding that interacts with the decoder and feature maps through multi-scale deformable attention, MOTR predicts target positions and classes in each frame. Notably, it establishes implicit temporal associations without the need for explicit data association or post-processing, seamlessly integrating detection and tracking to improve accuracy and robustness. The introduction of Track Query Aggregation with Learned Attention (TALA) is a key innovation, which models long-term temporal relations using a query memory bank and multi-head attention. Additionally, MOTR introduces innovative concepts such as Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to address target conflicts and enhance the integration of temporal information. These design elements collectively position MOTR as a significant milestone in multi-object tracking, elevating both its performance and applicability.
@@
MOTRv2 [47] enhances MOTR’s detection performance by incorporating the pre-trained object detector YOLOX [11] for detection priors. This involves anchoring object queries and utilizing YOLOX-generated proposal boxes as anchors to initialize the object queries. This strategy not only improves detection but also mitigates conflicts between detection and association tasks while preserving query propagation characteristics.
MeMOTR [10] introduces track queries, modeling target trajectories throughout the video and predicting them iteratively in each frame. It implements a tracklet-aware label assignment method to achieve one-to-one assignment between track queries and target trajectories, considering newborn object queries.
MOTRv3 [40] and CO-MOT [37] share a common focus on addressing unfair label assignment between detection and tracking queries in MOTR, aiming to balance training samples and improve end-to-end multi-object tracking performance. MOTRv3 employs a release-fetch supervision strategy for balanced label assignment, while CO-MOT uses a coopetition label assignment strategy to increase positive samples for detection queries and introduces "shadows" to enrich query quantity and diversity.
2.2 YOLO
YOLO[25] has transformed object detection into a regression model, achieving efficient target detection. It partitions the image into grids, with each grid responsible for predicting the target box and class information within its corresponding region. The object detection task is treated as a regression problem for coordinates and dimensions. This unique approach simplifies the detection process, enabling YOLO to simultaneously process multiple targets in a single forward pass, thus achieving outstanding speed and efficiency. The concept of this regression model provides a solid foundation for real-time object detection.
Due to its relatively high efficiency and accuracy, the SORT[3] initially applied YOLO in the field of MOT. However, with the proliferation of such models, YOLO has become widely used in the MOT domain. Nevertheless, this approach still introduces additional processing layers, potentially leading to some errors. Therefore, in this paper, we construct an end-to-end tracking model based on the YOLO framework.
2.3 RT-DETR
DETR [50] revolutionizes object detection by introducing the Transformer architecture, transforming the task into a multi-class classification problem. It excels in precise object localization and modeling object relationships, making it particularly suitable for applications like multi-object tracking. Despite its effectiveness with limited annotated data, DETR comes with the drawback of high computational complexity, demanding significant computational resources.
RT-DETR [21] addresses DETR’s computational challenges, presenting the first real-time usable DETR model. It tackles complexity with an efficient hybrid encoder that manages multi-scale features through smart intra-scale interactions and cross-scale fusion decoupling. This significantly reduces computational costs, achieving real-time object detection.
RT-DETR’s decoder module incorporates deformable attention, aligning with attention mechanisms in various DETR models [50, 36, 51]. By dynamically generating attention sampling points, it adapts to diverse feature distributions and target shapes, enhancing adaptability. Unlike some DETR models [36, 43, 8], RT-DETR skips the use of a reconstruction network, relying on the decoder’s output for final predictions. This simplifies the architecture while maintaining efficiency and performance.
3 MO-YOLO
@@
Figure 2: The overall structure of MO-YOLO consists of an architecture that combines YOLOv8’s backbone and neck, facilitating the extraction of image features for each frame. By merging the detect queries ($q_{d}$) and track queries ($q_{tr}$), the decoder (Dec) processes them to generate hidden states. When the model is trained using the TALA strategy, the TBSP can be skipped.
3.1 MO-YOLO Architecture
@@
MO-YOLO’s overall architecture is illustrated in the provided Fig.2. In this setup, a video sequence is input to YOLOv8’s backbone, neck to extract features. For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the RT-DETR decoder. The learning pattern of $q_{d}$ is similar to RT-DETR[21].
@@ 3.2 Tracking Box Selection Process
When training a detection model using the detection method and subsequently employing an end-to-end tracking approach, we observed a gradual increase in the number of tracking boxes for the same object over time. Upon examination, this phenomenon arises from the fact that each newly generated detect query is consistently assigned to an already tracked object. This is due to the nature of DETR, where a detect query can potentially be assigned to any object in the image, as label assignment is determined through binary matching of all detect queries and ground truth.
In MOTR[42], a solution to this issue is introduced through the utilization of TALA. This strategy gradually suppresses the generation of detect queries similar to track queries during the training process, effectively addressing the identified problem. However, this process increases the number of parameters that the network needs to automatically adjust, resulting in a prolonged convergence time for the model. Therefore, this paper proposes a simple yet effective approach named tracking box selection process (TBSP). Experimental results demonstrate the efficacy of TBSP, revealing that this method allows the model to achieve comparable performance to fully trained models even without extensive training. Additionally, the simplicity of TBSP ensures that it introduces minimal additional computational burden.
Figure 3: Refinement process for MO-YOLO on partially trained track queries. When MO-YOLO undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized. At that point, the track query generation method of MO-YOLO aligns with MOTR, as illustrated in the Fig. 1.
@@
Figure 4: The specific process of TBSP is represented by a formula. The specific meaning of the formulas in the figure refers to Section 3.2.
@@
The feasibility of TBSP is illustrated in the Fig. 4. Incidentally, although TBSP makes MO-YOLO appear not to be a truly end-to-end network, after thorough TALA training, this step can be omitted. Simultaneously, to better illustrate the performance of our model compared to MOTR in CO-MOT[38] and MOTRv3[40], we have not adopted the labeling strategy used in CO-MOT and MOTRv3.
3.3 Decoder Embeddings and Query Positions
Differentiating itself from MOTR, MO-YOLO’s decoder employs a methodology akin to RT-DETR for its embeddings. In contrast to MOTR, MO-YOLO’s decoder input doesn’t originate from a distinct encoder; rather, it exploits RT-DETR’s IoU-aware query selection mechanism. Building upon prior works such as [35, 43, 50, 16], this process strategically selects high-quality image features as the initial object embeddings for the decoder, emphasizing IoU scores.
It’s essential to clarify that in this context, the term "embeddings" specifically pertains to the initial embeddings of detection queries. The subsequent embeddings of track queries undergo generation through a TAN network, aligning with MOTR’s established process.
To expedite model convergence, we implement a Query Positions generation strategy akin to CO-MOT. Drawing inspiration from DAB-DETR [19], we employ a fixed transformation to convert reference points into Query Positions. While experimentation with the RT-DETR approach involving a Multilayer Perceptron (MLP) to transform reference points shows no significant performance discrepancy, it is important to note that the use of MLP substantially increases training time. By the way, it is worth mentioning that this process differs significantly from MOTR. In MOTR, Query Positions are generated using a set of learnable parameters, whereas in MO-YOLO, they are generated using reference points.
@@ 3.4 Loss and Training Methods
Due to the unique training mode adopted by MOTR, its batch size in a single GPU is constrained to 1. While this training approach enables the model to effectively learn the temporal information of the targets, it concurrently requires training the network to extract features of target appearances. To expedite the model training speed, a three-stage training strategy is employed in this study. In the first stage, the emphasis is placed on training the network to extract features of target appearances; the second stage focuses on learning temporal information; and the final stage comprehensively enhances the model’s performance by integrating the learning of temporal information and appearance features. This strategy is implemented to overcome the constraints posed by the unique training mode of MOTR, aiming to achieve more efficient model training.
@@
$\mathcal{L}$ is the loss of single frame, which is similar to the detection loss in DETR, and it is shown in Eq.3.
$\mathcal{L}\left(\left.\widehat{Y}_{t}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}$ (3)
where $\hat{\omega}_{tr}^{t-1}$ represents the multi-frame predictions,$\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the loss calculated over the entire video sequence, using ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$ $\mathcal{L}_{cls}$ denotes the focal loss[17], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[26], $\lambda_{cls},\lambda_{l1},\lambda_{giou}$ are corresponding weight coefficients.
@@
The second stage of training is conducted using MOTR’s Collective Average Loss. During this phase, only MOTR’s TAN and our TBSP are utilized. The loss functions used in this stage and the third stage are as Eq.4. Simultaneously, it is important to note that, at this juncture, the determination of the newly generated track queries relies solely on the scores of object queries and whether they have been filtered out by TBSP, as more advanced strategies like TALA have not been employed.
@@
The third stage of training involves the utilization of TALA policy during the training phase, while simultaneously deactivating the TBSP. Experimental evidence indicates that if the training epochs during this stage are sufficient, TBSP can also be disabled during the inference phase without significantly compromising performance. This finding underscores the robustness of the model, suggesting that once adequately trained in the third stage, the TBSP policy can be turned off during inference with negligible impact on overall performance.
In addition, it is important to note that while the first and second stages of MO-YOLO share similarities in generating a new track query from the object query, discrepancies may arise in the third stage. Tracking objects with the track query does not involve additional techniques, and the process across the three stages of MO-YOLO closely resembles that of MOTR.

### v4

# DecoderTracker: Decoder-Only Method for Multiple-Object Tracking

@@ Abstract
Decoder-only models, such as GPT[1], have demonstrated superior performance in many areas compared to traditional encoder-decoder structure transformer models. Over the years, end-to-end models based on the traditional transformer structure, like MOTR, have achieved remarkable performance in multi-object tracking. However, the significant computational resource consumption of these models leads to less friendly inference speeds and training times. To address these issues, this paper attempts to construct a lightweight Decoder-only model: DecoderTracker for end-to-end multi-object tracking. Specifically, drawing on some real-time detection models, we have developed an image feature extraction network which can efficiently extract features from images to replace the encoder structure. In addition to minor innovations in the network, we analyze the potential reasons for the slow training of MOTR-like models and propose an effective training strategy to mitigate the issue of prolonged training times. On the DanceTrack dataset, without any bells and whistles, DecoderTracker’s tracking performance slightly surpasses that of MOTR, with approximately twice the inference speed. Furthermore, DecoderTracker requires significantly less training time compared to MOTR.
@@ 1 Introduction
Multi-Object Tracking (MOT) is a critical task in the field of video analysis, primarily aimed at inferring and predicting the motion trajectories of objects or instances across a sequence of continuous images. The success of transformer[2] models in various domains in recent years is well-documented, and recently, transformer-based end-to-end MOT models like MOTR[3] have been introduced in the MOT domain. These models offer advantages over the widely used tracking-by-detection methods[4, 5, 6, 7], such as eliminating the need for manual design or selection of features. Additionally, some enhanced methods[8, 9, 10] of MOTR have demonstrated the potential of this model class in terms of tracking accuracy.
However, the inference speed of these models is often suboptimal, and their RNN[11] like linear inference structure makes them less GPU friendly during training.
[…]
At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the model to engage in less stringent learning might also be effective. Is there a possibility that this process could yield unexpected benefits? Based on this idea, we proposed the Tracking Box Selection Process (TBSP).
Although the methods we propose have not yet enabled the inference speed of MOTR-like models to achieve real-time performance, we believe that this paper could provide valuable insights for future researchers.
In summary, the contributions of this paper are as follows:
@@
    The primary contribution involves the development of a novel end-to-end tracking network, DecoderTracker, which is a decoder-only method. Compared to MOTR, this network achieves faster inference speeds and improved tracking performance.
@@
    The third contribution introduces TBSP, a strategy that can serve as a preliminary weakly supervised training approach for end-to-end MOT model training.
@@ 2 Related work
Figure 1: MOTR series’ Streamline the process of updating detect queries and track queries under typical MOT scenarios. The track query set, initially empty, is dynamically adjusted in length. Detect queries play a role in identifying new objects.
@@ 2.1 MOTR Series
Unlike tracking-by-detection models exemplified by SORT[4, 14, 5], and previous transformer-based MOT models[15, 16], MOTR[3] is the first model to achieve end-to-end multi-object tracking. It extends the object query from DETR[17] into a track query, thereby enabling tracking capabilities. It uses a learnable positional embedding and multi-scale deformable attention to predict object positions and classes in frames, eliminating the need for explicit data association or post-processing. Key innovations include the Track Query Aggregation with Learned Attention (TALA), which uses a query memory bank and multi-head attention for modeling long-term temporal relationships, and the introduction of Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to resolve target conflicts and improve temporal information integration. These features establish MOTR as a significant advancement in multi-object tracking technology.
@@
MOTRv2[18] enhances the detection capabilities of the original MOTR by integrating the pre-trained object detector YOLOX[19]. This integration involves using YOLOX generated proposal boxes as anchors to initialize object queries, a method that not only bolsters detection accuracy but also helps to alleviate conflicts between detection and association tasks, all the while maintaining the query propagation features inherent to MOTR.
Similarly, MeMOT[10] innovates by incorporating track queries that model and predict target trajectories iteratively across video frames. This approach utilizes a tracklet-aware label assignment method to ensure a one-to-one correspondence between track queries and actual target trajectories, addressing challenges with newly introduced object queries.
Moreover, MOTRv3[9] and CO-MOT[20], as detailed in their respective papers, focus on rectifying the imbalance in label assignment between detection and tracking queries within the MOTR framework. MOTRv3, adopts a release-fetch supervision strategy to facilitate balanced label assignments, enhancing overall MOT performance. In contrast, CO-MOT, employs a coopetition label assignment strategy that not only increases the number of positive samples for detection queries but also introduces "shadows" to enhance the diversity and quantity of queries, further refining the MOT capabilities.
2.2 Real Time Detection Method
When discussing real-time detection models, most people initially think of YOLO[21], which transformed the problem of object detection into a regression model, achieving efficient object detection. The model divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions. The most advanced versions in the YOLO series currently are YOLOv8.2[22] and YOLOv9[23].
Recently, the real-time detection model based on DETR, known as RT-DETR[24], has also demonstrated its extraordinary performance. RT-DETR achieves optimized inference speeds by designing an efficient encoder and simplifying the model structure.
Although we acknowledge the efficiency of RT-DETR’s hybrid encoder, we believe that the encoder structure might be unnecessary. Therefore, inspired by the YOLO series[25, 22, 26], we designed a feature extraction network.
3 DecoderTracker
@@
Figure 2: The overall structure of DecoderTracker. By merging the detect queries ($q_{d}$) and track queries ($q_{tr}$), the decoder (Dec) processes them to generate hidden states. When the model is trained using the TALA strategy, the TBSP can be skipped.
Figure 3: The structure of IMFENet. It bears certain similarities to the structure of YOLOv8’s backbone and neck, yet there are distinct differences. For detailed comparisons, please refer to the appendix.
3.1 DecoderTracker Architecture
@@
DecoderTracker’s overall architecture is illustrated in the provided Fig.2. In this setup, a video sequence is input to Image Feature Extraction Network(IMFENet) to extract features. For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the decoder.
@@
The structure of IMFENet is primarily inspired by YOLOv8, which replaces the backbone and encoder structures of the original MOTR for image feature extraction. The reason for not adopting the latest YOLOv9 is due to its slower training speed compared to YOLOv8. Through IMFENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder.
@@ 3.2 Tracking Box Selection Process
Figure 4: The specific process of TBSP is represented by a formula. The specific meaning of the formulas in the figure refers to Section 3.2
The initial design of the TBSP was intended merely to equip a detection network with basic tracking functionality. We observed that when a detection network directly employs the MOTR[3] method to update track queries, the tracking boxes for the same object tend to increase over time. This accumulation was found to be due to each newly generated detect query invariably being assigned to an already tracked object. In the DETR, a detect query can be assigned to any object in the image, as label assignment is determined through bipartite matching between all detection queries and ground truth data. MOTR addresses this by utilizing tracklet-aware label assignment (TALA), which suppressively modulates the generation of detect queries similar to track queries during training, thus mitigating this issue.
Consequently, we designed a simple filtering mechanism, TBSP, which enables the detection network to possess some tracking capabilities without additional training. However, given the network’s inherently weak tracking ability and the lack of robustness in TBSP, the tracking performance at this stage is quite limited. Nevertheless, we also discovered that using TBSP as a preprocessing step for TALA, and training the network in this fashion, can lead to some unexpected benefits.
Compared to TALA, TBSP can be considered a form of self-supervised or weak supervision during training. This is because TBSP does not forcibly assign a specific query to correspond to an object; instead, it allows the model to optimize itself. For a discussion on why TBSP is effective, please refer to the appendix.
@@
The feasibility of TBSP is illustrated in the Fig. 4. Incidentally, although TBSP makes DecoderTracker appear not to be a truly end-to-end network, after thorough TALA training, this step can be omitted. Simultaneously, to better illustrate the performance of our model compared to MOTR, we have not adopted the labeling strategy used in CO-MOT[8] and MOTRv3[9].
@@ 3.3 Loss and Training Methods
Due to MOTR’s similarity to RNN in its inference approach, its batch size is limited to 1. This often results in suboptimal utilization of GPU resources during training. Additionally, MOTR learns both image and motion features simultaneously. To expedite model training, this paper adopts a three-stage training strategy. In the first stage, emphasis is placed on training the network to extract target image features efficiently. The second stage involves training the model on weak tracking aspects using TBSP as a pretraining step for the third stage. Lastly, the third stage employs the TALA strategy to comprehensively enhance the model’s tracking performance.
@@
$\mathcal{L}$ is the loss of single frame. Notably, during this stage, we employed the IoU-aware Query Selection from RT-DETR, so the loss function utilized is from RT-DETR. Unlike the original DETR, this stage introduces a loss termed $\mathcal{L}_{enc}$. This loss involves treating the initialized queries as boxes for the purpose of loss computation. The specifics are outlined below:
$\mathcal{L}_{enc}(\hat{y},y)&=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},\hat{b},y,b)\\ &=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},c,IoU)\end{aligned}$ (3)
where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}={\hat{c},\hat{b}}$ and $y={c,b}$, $c$ and $b$ represent categories and bounding boxes, respectively. Finally, the overall loss is as follows:
$\mathcal{L}\left(\left.\widehat{Y_{t}}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}+\lambda_{enc}\mathcal{L}_{enc}$ (4)
where $\hat{\omega}_{tr}^{t-1}$ represents the multi-frame predictions,$\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the loss calculated over the entire video sequence, using ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$ $\mathcal{L}_{cls}$ denotes the focal loss[27], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[28], $\lambda_{cls},\lambda_{l1},\lambda_{giou},\lambda_{enc}$ are corresponding weight coefficients.
@@
The second stage of training is conducted using MOTR’s Collective Average Loss. During this phase, only MOTR’s TAN and our TBSP are utilized. The loss functions used in this stage and the third stage are as Eq.5.
@@
At this stage, the computation of $\mathcal{L}_{enc}$ also differs from that in the first phase, where the newly generated detection queries are calculated only in conjunction with newly appeared objects. Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to it in second stage.
The third stage of training involves the utilization of TALA during the training phase, while simultaneously deactivating the TBSP. Experimental evidence indicates that if the training epochs during this stage are sufficient, TBSP can also be disabled during the inference phase without significantly compromising performance. This finding underscores the robustness of the model, suggesting that once adequately trained in the third stage, TBSP can be turned off during inference with negligible impact on overall performance.